OpenStudy (quien125):

Line m has no y-intercept, and its x-intercept is (3, 0). Line n has no x-intercept, and its y-intercept is (0, -4). what is the equation for line m and n?

OpenStudy (3mar):

May I help?

OpenStudy (quien125):

go for it

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (quien125):

yepp

OpenStudy (3mar):

But I did not solve for anything yet!

OpenStudy (3mar):

Anyway...

OpenStudy (quien125):

alright

OpenStudy (3mar):

Can you describe the status of the first line?

OpenStudy (3mar):



OpenStudy (quien125):

what would be the equatoin in y=mx+b

OpenStudy (3mar):

Good. Nice question. For the form you wrote, m stands for "slope" and b stands for the intercept with y-axis. Consequently, no intersection with y, then?

OpenStudy (quien125):

nope so it would be y=x+3 for line m and line n is y=-4x right?

OpenStudy (3mar):

One moment, please. why you say "y=x+3"?

OpenStudy (3mar):

Then the slope with be 1, and this is not correct because the slope of this line that intercept x-axis at (3,0) and has no intercept with y-axis (i.e parallel to y-axis) is undefined. Do you follow?

OpenStudy (3mar):

The equation of line m is so simple : just x=3 as this value does not change whatever y-value changes

OpenStudy (3mar):

@quien125

OpenStudy (quien125):

ya i did it it in slope intercept form

OpenStudy (3mar):

Where were you? I was waiting!

OpenStudy (quien125):

sorry went to the store for my sisters candles for her bday cake

OpenStudy (3mar):

Ok

OpenStudy (3mar):

So we solved for m, right?

OpenStudy (quien125):

yes and n was y=-4x

OpenStudy (quien125):

right in y intercept form

OpenStudy (3mar):

Why y=-4"x"? what if x changes, does y change?

OpenStudy (quien125):

it doesn't. i just need the equation in yintercept form for both

OpenStudy (3mar):

So it will be just y=-4. Is not it?

OpenStudy (quien125):

but x would need to be included to make it a y=mx+b equation

OpenStudy (quien125):

just after -4

OpenStudy (3mar):

Try this.

OpenStudy (3mar):

You say "but x would need to be included just after -4"!! Try to put it and see if it satisfy or not.

OpenStudy (quien125):

alright

OpenStudy (3mar):

Of course it will not. These lines are parallel lines to the x and y-axes, one of them (parallel to x-axis) has zero slope (m=0) and the other (parallel to y-axis) has undefined slope (m=1/0)

OpenStudy (quien125):

oh i see thank you. i see where i went wrong.

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (quien125):

thank you for helping!!

OpenStudy (3mar):

You are welcome!

563blackghost (563blackghost):

What site is that? The `2.png` pic.

563blackghost (563blackghost):

@3mar . Sorry didn't specify who I am talking to.

OpenStudy (3mar):

Never mind.

OpenStudy (3mar):

It is the coolest platform I have ever seen for graphing and other cool things https://www.calorful.com/graphing/calculator Feel free to look around and test your graphs and equations.

OpenStudy (3mar):

@quien125 got the last pic?

563blackghost (563blackghost):

Cool xD Thank you! @3mar

OpenStudy (3mar):

Not at all!